One vertical angle is a number plus 12. The other is 40 less than 3x the same number. What is the number??
Answer by RAY100(1637)

Vertical angles are equal
.
(n+12) = 3n-40
.
52 = 2n
.
26 =n
.
One angle = (n+12) = 38
.
Other angle = 3*26 - 40 = 78-40 = 38,,,,,ok
.